COHERENCE-CONVERTED POPULATION TRANSFER FTMW-IR DOUBLE RESONANCE SPEC-TROSCOPY OF CH3OD IN THE C-H STRETCH REGION

\underline{SYLVESTRE TWAGIRAYEZU}, DAVID S. PERRY, Department of Chemistry, The University of Akron, Akron OH 44325; JUSTIN L. NEILL, MATT T. MUCKLE, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904.

Coherence-converted population transfer microwave-infrared double resonance spectroscopy is employed to record the rotationally state-selected infrared spectra of jet-cooled CH_{3}OD in the C-H stretch region (2750−3020 cm^{−1}). The observed infrared spectra result from the E-species microwave transitions (1_{0} ← 1_{−1} at 18.957 GHz, 2_{0} ← 2_{−1} at 18.991 GHz, and 3_{0} ← 3_{−1} at 19.005 GHz). The present spectra of CH_{3}OD contain 17 interacting vibrational bands (J′ = 0). In additional to the three C-H stretch fundamentals (ν_{3}:2841.7 cm^{−1}, ν_{9} : 2954.4cm^{−1} and ν_{2} : 2998.9cm^{−1}), 14 additional band origins are found in the region of the binary combinations of the CH bends (2890−2950 cm^{−1}). Although the A-species was inaccessible in the present work, the pattern of E-species reduced energies suggests that the torsional tunneling splittings of ν_{3} and ν_{9} are normal, whereas ν_{2} is inverted. The number and distribution of the observed vibrational bands support a stepwise coupling scheme in which the CH stretch bright state couples first to the binary C-H bend combinations, and then to all of the higher order vibrational combinations. A time-dependent interpretation in the asymmetric region indicates a fast (170 fs) initial decay of the bright state.